# KEK-TH-XXXX

### v1

@@ 4.1 Way to exclude the memory burden effect
A way to avoid this conclusion is to dilute the overproduced memory-burdened PBH dark matter though this also dilutes the baryon asymmetry of the Universe if the generation of the baryon asymmetry took place earlier. In addition, the (incomplete) evaporation of the PBHs produces entropy and dilutes the baryon asymmetry further. The PBHs initially heavier than about $10^{9}\,\mathrm{g}$ are tightly constrained by the BBN. If one assumes that the baryon asymmetry of the Universe was produced sufficiently early (before PBH evaporation), the observed baryon-to-photon ratio puts an upper bound on the initial PBH abundance $\gamma\beta=(\rho_{\text{PBH}}/\rho_{\text{total}})|_{\text{ini}}$: $\beta^{\prime}\lesssim 10^{-5}\left(\frac{10^{9}\,\mathrm{g}}{M_{\text{PBH,ini}}}\right)$ for $M_{\text{PBH,ini}}<10^{9}\,\mathrm{g}$, where $\beta^{\prime}\equiv\gamma^{3/2}\left(\frac{106.75}{g_{*,\text{ini}}}\right)^{1/4}\left(\frac{0.67}{h}\right)^{2}\beta$ (see Ref. [4] and references therein). [^7] The definition of $\beta$ in this paper is different from that in Ref. [4]: $\gamma\beta_{\text{here}}=\beta_{\text{there}}$, where $\beta_{\text{here}}$ is $\beta$ in this paper (the formation probability) and $\beta_{\text{there}}$ is $\beta$ defined in Ref. [4] (the initial abundance). Note that this bound can be evaded if the baryon asymmetry of the Universe is produced after the PBH evaporation and before the BBN.
@@
4.2 Ways to confirm the memory burden effect
@@
Optimistically, an indirect hint of PBH dark matter may be already found. In Ref. [106], it was found that the rate of the fast radio bursts (FRBs) increases in the presence of PBHs. The FRB rate asymptotes to a value comparable with observations [107, 108] for PBHs lighter than about $10^{22}\,\mathrm{g}$ [106]. If we can exclude the conventional PBH mass range for the entirety of dark matter and other astrophysical explanations, we can argue that the memory-burdened PBHs are responsible for the observed FRB rate.
Next, we discuss the GWs originating from mergers of the memory-burdened PBH dark matter. This has been well studied in the context of explaining the LIGO/Virgo GW signals by PBH binary mergers. PBHs can form a binary in the early Universe [109, 110, 111, 63] as well as in the late Universe [112, 113]. For $30\,M_{\odot}$ PBHs, the former mechanism is more effective [111, 63]. The merger rate per comoving volume approximately scales as $M_{\text{PBH}}^{-32/37}$ and $M_{\text{PBH}}^{0}$ for the early and late formation mechanisms, respectively, so the early-time binary formation mechanism is dominant also for PBHs lighter than $30\,M_{\odot}$. The spectrum of the stochastic GW background from the superposition of mergers of light PBH binaries was discussed in the literature [114, 115, 50, 116, 66]. [^8] The same calculation applies to the supermassive black hole binaries if the environmental effect is negligible. If the evidence [117, 118, 119, 120] of the GWs around nanohertz found by the worldwide pulsar timing array collaborations originates from the mergers of the supermassive black hole binaries, the environmental effect must be significant. See Refs. [121, 122] and references therein. It is obtained as follows:
@@
where $\rho_{\text{c}}$ is the critical density, $z_{\text{sup}}$ is the upper bound of the redshift integral, $R$ is the merger rate of PBH binaries per comoving volume, $f_{\text{s}}=(1+z)f$ is the frequency in the source frame, and $\mathrm{d}E/\mathrm{d}f_{\text{s}}$ is the GW spectrum from a merger in the source frame [123, 124]. The expressions of the upper bound $z_{\text{sup}}$, the merger rate $R$, and the energy spectrum $\mathrm{d}E/\mathrm{d}f_{\text{s}}$ are given in Appendix A. The spectrum of $\Omega_{\text{GW}}^{\text{(merger)}}$ is shown in Fig. 3. The spectral shape is the same as the previous work in the literature. It follows the power-law $\Omega_{\text{GW}}^{\text{(merger)}}\propto f^{2/3}$ [125, 123, 124] below the peak frequency. The peak frequency is $f_{\text{peak}}=2\times 10^{27}\,(M_{\text{PBH,ini}}/(10^{10}\,\mathrm{g}))^{-1}\,\mathrm{Hz}$, and the intensity at the peak is $\Omega_{\text{GW}}^{\text{(merger)}}(f_{\text{peak}})h^{2}=6\times 10^{-9}(M_{\text{PBH,ini}}/(10^{10}\,\mathrm{g}))^{5/37}$. Currently, there are no known methods to detect such extremely high-frequency GWs. For recent developments toward the detection of high-frequency GWs, see Refs. [126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140].
[…]
[width=0.6]GWspectrum_merger3.pdf
Figure 3: The spectra of the GWs from mergers of memory-burdened PBH binaries. From left to right, $M_{\text{PBH,ini}}/\mathrm{g}=10^{10}$ (the blue solid line), $10^{5}$ (the orange dashed line), and 1 (the green dot-dashed line) with $M_{\text{PBH}}=M_{\text{PBH,ini}}/2$. The monochromatic mass function is assumed for simplicity. The same sensitivity curves as in Fig. 1 are also shown at the top left corner.
@@ Acknowledgment
K. K. and T. T. thank Sai Wang for discussions on the GWs from mergers of PBHs. This work was in part supported by MEXT KAKENHI Grants No. JP23KF0289, No. JP24H01825, No. JP24K07027 (K. K.), and No. 24H02244 (T. T. Y.). T. T. Y. was supported also by the Natural Science Foundation of China (NSFC) under Grant No. 12175134 as well as by World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan. T. T.’s work was supported by the RIKEN TRIP initiative (RIKEN Quantum).
@@
In this appendix, we summarize the formulas to calculate the spectrum of the GWs from mergers of PBH binaries. We follow Ref. [63] as well as Appendixes B and C of Ref. [115]. See also Appendix B of Ref. [50]. For simplicity, we assume a monochromatic PBH mass spectrum and the homogeneous spatial distribution of the PBHs. Also, we use the late-time PBH mass (in the memory burden regime) and do not take into account the time dependence of the PBH mass. This is justified since light PBHs quickly enter the memory burden regime.
@@
The merger rate per comoving volume is [63]
@@
where $M_{\text{c}}=2^{-1/3}M_{\text{PBH}}$ is the chirp mass, $f_{1}=0.4499M_{\text{PBH}}^{-1}$, $f_{2}=1.026M_{\text{PBH}}^{-1}$, $f_{3}=1.401M_{\text{PBH}}^{-1}$ and $\sigma=0.2381M_{\text{PBH}}^{-1}$, where we have simplified the expressions assuming the monochromatic spectrum of the PBHs. Remember that we use $8\pi G=1$ rather than $G=1$.
@@
Figure 4: The contour plot of $\log_{10}\text{SNR}$ for 1-year observation on the plane of $\log_{10}(M_{\text{PBH,ini}}/\text{g})$ and $\log_{10}\Delta$. The border between the pinkish (low SNR) and greenish (high SNR) regions corresponds to $\text{SNR}=1$ and the $\log_{10}\text{SNR}$ changes by 11 for the adjacent contours. The corresponding figure for CE is shown on the right panel of Fig. 2

### v2

@@ 4.1 Way to exclude the memory burden effect
There is a caveat on this program. If one observes only the IR tail of the GW spectrum, the information of the amplitude and of the peak frequency, or equivalently, the PBH abundance and PBH mass, are degenerate because a smaller amplitude with a lower peak frequency has a similar signal with a larger amplitude with a higher peak frequency. To exclude the memory burden effect, one first needs to measure the GW spectrum in a sufficiently broad range so that the degeneracy is resolved.
Suppose that one has excluded the memory burden effect by observing the induced GWs stronger than expected. In this case, if the PBHs were so abundant that they became the dominant component of the Universe, their evaporation produce the hot big-bang Universe associated with gravitons, which contribute to the effective number of species of neutrinos $N_{\text{eff}}$. The excess can be detected by future observations such as CMB-S4 [114] if the PBHs acquire substantial spin, e.g., by merger processes [115, 116].
On the other hand, a way to avoid exclusion of the memory burden effect is to dilute the overproduced memory-burdened PBH dark matter by entropy production due, e.g., to heavy particle decay (see, e.g., Refs. [117, 118, 119, 120, 121]) or to thermal inflation [122], though this also dilutes the baryon asymmetry of the Universe if the generation of the baryon asymmetry took place earlier. In addition, the (incomplete) evaporation of the PBHs produces entropy and dilutes the baryon asymmetry further. The PBHs initially heavier than about $10^{9}\,\mathrm{g}$ are tightly constrained by the BBN. If one assumes that the baryon asymmetry of the Universe was produced sufficiently early (before PBH evaporation), the observed baryon-to-photon ratio puts an upper bound on the initial PBH abundance $\gamma\beta=(\rho_{\text{PBH}}/\rho_{\text{total}})|_{\text{ini}}$: $\beta^{\prime}\lesssim 10^{-5}\left(\frac{10^{9}\,\mathrm{g}}{M_{\text{PBH,ini}}}\right)$ for $M_{\text{PBH,ini}}<10^{9}\,\mathrm{g}$, where $\beta^{\prime}\equiv\gamma^{3/2}\left(\frac{106.75}{g_{*,\text{ini}}}\right)^{1/4}\left(\frac{0.67}{h}\right)^{2}\beta$ (see Ref. [4] and references therein).999 The definition of $\beta$ in this paper is different from that in Ref. [4]: $\gamma\beta_{\text{here}}=\beta_{\text{there}}$, where $\beta_{\text{here}}$ is $\beta$ in this paper (the formation probability) and $\beta_{\text{there}}$ is $\beta$ defined in Ref. [4] (the initial abundance). Note that this bound can be evaded if the baryon asymmetry of the Universe is produced after the PBH evaporation and before the BBN.
@@
4.2 Way to confirm the memory burden effect
@@
Let us discuss the GWs originating from mergers of the memory-burdened PBH dark matter. This has been well studied in the context of explaining the LIGO/Virgo GW signals by PBH binary mergers. PBHs can form a binary in the early Universe [123, 124, 125, 68] as well as in the late Universe [126, 127]. For $30\,M_{\odot}$ PBHs, the former mechanism is more effective [125, 68]. The merger rate per comoving volume approximately scales as $M_{\text{PBH}}^{-32/37}$ and $M_{\text{PBH}}^{0}$ for the early and late formation mechanisms, respectively, so the early-time binary formation mechanism is dominant also for PBHs lighter than $30\,M_{\odot}$. The spectrum of the stochastic GW background from the superposition of mergers of light PBH binaries was discussed in the literature [128, 129, 55, 130, 71].101010 The same calculation applies to the supermassive black hole binaries if the environmental effect is negligible. If the evidence [131, 132, 133, 134] of the GWs around nanohertz found by the worldwide pulsar timing array collaborations originates from the mergers of the supermassive black hole binaries, the environmental effect must be significant. See Refs. [135, 136] and references therein. It is obtained as follows:
@@
where $\rho_{\text{c}}$ is the critical density, $z_{\text{sup}}$ is the upper bound of the redshift integral, $R$ is the merger rate of PBH binaries per comoving volume, $f_{\text{s}}=(1+z)f$ is the frequency in the source frame, and $\mathrm{d}E/\mathrm{d}f_{\text{s}}$ is the GW spectrum from a merger in the source frame [137, 138]. The expressions of the upper bound $z_{\text{sup}}$, the merger rate $R$, and the energy spectrum $\mathrm{d}E/\mathrm{d}f_{\text{s}}$ are given in Appendix A. For the calculation of the merger rate, we follow Ref. [68] and neglect the effects of PBH cloud, which may disrupt the binary, once the binary is formed. See Ref. [139] for the effects of the surrounding PBHs, which may significantly affect the merger rate.
The spectrum of $\Omega_{\text{GW}}^{\text{(merger)}}$ is shown in Fig. 4. The spectral shape is the same as the previous work in the literature. It follows the power-law $\Omega_{\text{GW}}^{\text{(merger)}}\propto f^{2/3}$ [140, 137, 138] below the peak frequency. The peak frequency is $f_{\text{peak}}=2\times 10^{27}\,(M_{\text{PBH,ini}}/(10^{10}\,\mathrm{g}))^{-1}\,\mathrm{Hz}$, and the intensity at the peak is $\Omega_{\text{GW}}^{\text{(merger)}}(f_{\text{peak}})h^{2}=6\times 10^{-9}(M_{\text{PBH,ini}}/(10^{10}\,\mathrm{g}))^{5/37}$. Currently, there are no known methods to detect such extremely high-frequency GWs. For recent developments toward the detection of high-frequency GWs, see Refs. [141, 142, 143, 144, 145, 83, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155].
Figure 4: The spectra of the GWs from mergers of memory-burdened PBH binaries. The PBHs are assumed to comprise the totality of dark matter. From left to right, $M_{\text{PBH,ini}}/\mathrm{g}=10^{10}$ (the blue solid line), $10^{5}$ (the orange dashed line), and 1 (the green dot-dashed line) with $M_{\text{PBH}}=M_{\text{PBH,ini}}/2$. The monochromatic mass function is assumed for simplicity. The same sensitivity curves as in Fig. 2 are also shown.
@@ 5 Conclusions
Note Added
Soon after our paper, Ref. [156] appeared on arXiv, which also discusses the induced GWs to constrain the PBH abundance in the context of the memory burden effect. There are some differences about the considered power spectra of curvature perturbations and the prescription to calculate the PBH abundance. While we focus more on the future prospects, they focus more on the current constraints from LIGO-Virgo O1-O3 runs. In this work, we additionally studied ways to either confirm or exclude the memory burden effect in Sec. 4.
@@ Acknowledgment
K.K. and T.T. thank Sai Wang for discussions on the GWs from mergers of PBHs. This work was in part supported by MEXT KAKENHI Grants No. JP23KF0289, No. JP24H01825, No. JP24K07027 (K.K.), and No. 24H02244 (T.T.Y.). T.T.Y. was supported also by the Natural Science Foundation of China (NSFC) under Grant No. 12175134 as well as by World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan. T.T.’s work was supported by the 34th Academic research grant FY 2024 (Natural Science) No. 9284 from DAIKO FOUNDATION and by the RIKEN TRIP initiative (RIKEN Quantum).
@@
In this appendix, we summarize the formulas to calculate the spectrum of the GWs from mergers of PBH binaries. We follow Ref. [68] as well as Appendixes B and C of Ref. [129]. See also Appendix B of Ref. [55]. For simplicity, we assume a monochromatic PBH mass spectrum and the homogeneous spatial distribution of the PBHs. Also, we use the late-time PBH mass (in the memory burden regime) and do not take into account the time dependence of the PBH mass. This can be justified since light PBHs quickly enter the memory burden regime and the evaporation in the memory burden regime is negligible for sufficiently large values of $n_{\text{MB}}$.
@@
Assuming the binary formation due to the torque from the nearest (third) PBH, the merger rate per comoving volume is [68]
@@
where $M_{\text{c}}=2^{-1/3}M_{\text{PBH}}$ is the chirp mass, $f_{1}=0.4499M_{\text{PBH}}^{-1}$, $f_{2}=1.026M_{\text{PBH}}^{-1}$, $f_{3}=1.401M_{\text{PBH}}^{-1}$ and $\sigma=0.2381M_{\text{PBH}}^{-1}$, where we have simplified the expressions assuming the monochromatic spectrum of the PBHs. Remember that we use $18\pi G=$ 1 rather than $G=1$.
As mentioned briefly in the main text, the comoving merger rate is modified by surrounding PBHs because a binary may be disrupted by the surrounding PBHs. This issue was studied in Ref. [139] with $N$-body simulations. Unfortunately, the analytic estimate in Ref. [139] is not directly applicable for the case $f_{\text{PBH}}=1$, but the authors of Ref. [139] provided a conservative estimate of the merger rate for $f_{\text{PBH}}=1$ and $M_{\text{PBH}}\approx 30M_{\odot}$. The merger rate derived as above and the conservative estimate in Ref. [139] deviate more for smaller PBH mass, which may suggest that the latter may be too conservative after the extrapolation of the PBH mass over more than 20 orders of magnitude into our parameter space $M_{\text{PBH}}\leq 10^{10}\,\mathrm{g}$. If such an extrapolation is valid, the merger rate is substantially suppressed, so are the resulting GWs.
@@
Figure 5: The contour plots of $\log_{10}\text{SNR}$ for 1-year observations of (a) ET, (b) BBO, (c) DECIGO, and (d) HLVK on the plane of $\log_{10}(M_{\text{PBH,ini}}/\text{g})$ and $\log_{10}\Delta$. The border between the pinkish (low SNR) and greenish (high SNR) regions corresponds to $\text{SNR}=1$ and the $\log_{10}\text{SNR}$ changes by $1$ for the adjacent contours. The corresponding figure for CE is shown on the right panel of Fig. 3